OpenStudy (anonymous):

f(x+h)-f(x)/h when f(x) = x^2 +1

OpenStudy (anonymous):

((x+h)^2+1 - (x^2+1))/h = (x^2+2xh+h^2-x^2-1)/h = (2xh+h^2)/h (other values cancel out) = (h(2x-h))/h = 2x- h (h's cancel out)

OpenStudy (anonymous):

hope that helps :)

OpenStudy (anonymous):

something went wrong with your signs

myininaya (myininaya):

\[f(x+h)=(x+h)^2+1=x^2+2xh+h^2+1\] \[-f(x)=-(x^2+1)=-x^2-1\] ----------------------------------add \[f(x+h)-f(x)=2xh+h^2\]

myininaya (myininaya):

\[\frac{f(x+h)-f(x)}{h}=\frac{2xh+h^2}{h}\]

myininaya (myininaya):

\[\frac{h(2x+h)}{h}\]

myininaya (myininaya):

Does anything cancel?